• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 21:35:47 on June15,2010

    Entry number 326815

    keyword=Lowered Current to 40 uA

    The target vacuum became less stable around 21:17 and was hovering about 6.5e10-5 level or worse. I had the beam current lowered to 40 uA and the vacuum has stabilized to about 3.5e10-5.